OpenStudy (anonymous):

CALCULUS HELP?!!! Please?… I really want to learn how to do this... Gas is escaping from a spherical balloon at the rate of 10 cubic feet per hour. At what rate is the radius changing when the volume is 36 pi?

OpenStudy (anonymous):

do you know the volume formula for a sphere?

OpenStudy (anonymous):

yes @dpaInc (4/3)pir^2

OpenStudy (anonymous):

well, not quite... \(\large V=\frac{4}{3}\pi r^3 \) taking the derivative with respect to time: \(\large V'=4\pi r^2\cdot r' \ before i explain further, do you understand how this was obtained?

OpenStudy (anonymous):

\(\large V'=4\pi r^2\cdot r' \) that's what it should read, sorry...

OpenStudy (anonymous):

yessss! I just saw that . Yes yes . Its understandable you took the derivative of that and thats how you got v'=4pir^2(r')

OpenStudy (anonymous):

ok.. V' represents the rate at which the volume is changing (given) r' is the rate at which the radius is changing (what we're asked to find) so \(\large r'=\frac{V'}{4\pi r^2} \), ok so far?

OpenStudy (anonymous):

mmmmm…? So you basically divided it. to get?….

OpenStudy (anonymous):

yes.. this is how we'll calclate r'... all we need is the values for V' and r... now you need to find r when the volume of the sphere is 36pi cubic feet.... do you know how to do that?

OpenStudy (anonymous):

noooo. I'll be honest. I don't know how to do that...

OpenStudy (anonymous):

ok... we need to find the radius, r, when the volume, V, is 36pi so we'll use the volume formula: \(\large V=\frac{4}{3} \pi r^3\) \(\large 36\pi=\frac{4}{3} \pi r^3\) now we solve for r... \(\large 36\cancel{\pi}=\frac{4}{3} \cancel{\pi} r^3\) \(\large 36\cdot \frac{3}{4}= r^3\) \(\large 27=r^3\) \(\large 3=r\) so radius is 3 when the volume is 36pi...

OpenStudy (anonymous):

one last thing, do you remember what V' = ???

OpenStudy (anonymous):

V' as in what it means? V' is prime?

OpenStudy (anonymous):

V' is the rate at which the volume is changing... it is given in the problem...

OpenStudy (anonymous):

V' is the rate at which gas is escaping (volume changing) the sphere.

OpenStudy (anonymous):

ok so how did 36pi take place of V?

OpenStudy (anonymous):

Ohhhhh Ok nvm that question …. I got it.

OpenStudy (anonymous):

since the gas is ESCAPING, this means the volume is changing -10 cubic feet per hour. so V' = -10. and we got r = 3 so now we'll put these values into the formula for r': \(\huge r'=\frac{V'}{4\pi r^2}=\frac{-10}{4\pi (3)^2} \)

OpenStudy (anonymous):

I was just about to ask what happens to the r'=v'/4pir^2

OpenStudy (anonymous):

-71?

OpenStudy (anonymous):

do you know what the units are for your answer?

OpenStudy (anonymous):

or -70.69?

OpenStudy (anonymous):

no... it's best to leave the answer in the form above (in terms of pi) if you want a decimal approximation then use your calculator....

OpenStudy (anonymous):

oh ok… Yes the units would be ft.

OpenStudy (anonymous):

ft per hour.

OpenStudy (anonymous):

yes exactly.

OpenStudy (anonymous):

thank you very muchhhhh! Im trying really hard to understand this.

OpenStudy (anonymous):

also notice that it's a negative answer which means the radius is decreasing (which makes sense because gas is escaping and the volume is getting smaller)

OpenStudy (anonymous):

yw...:)

OpenStudy (anonymous):

do you mind helping me with one more? @dpaInc